Operation Summary for November 22, 2010 - November 24, 2010 |
Monday Nov. 22 The day shift began with Operations monitoring Tevatron (TeV) store 8291 (184.64E30) and the Recycler stash (152.2E10), with the Antiproton Source (Pbar) stacking (26.4E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and the Meson FTBF T-992 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Instrumentation personnel rebooted the TeV sector C1 BPM house and recovered the 8 GeV line Multiwire system after it stopped responding. A tornado warning at 3:20 PM required operators to halt beam from the Proton Source, to clear the Main Control Room, and head to the shelters. Operations resumed Proton Source beam after the all-clear sounded. During the evening shift, at 4:35 PM, Operations reported that a lightning strike caused the D48 horizontal separator to spark and aborted store 8291. Operations put the TeV into shot setup and established store 8292 at 6:17 PM with an average initial luminosity (AIL) of 297.07E30. Operations disabled the T-992 beam at 6:30 PM. Operations reported several NuMI downtimes due to an exhaust stack fan (SR3FNI) that tripped and pulled the beam permit. Investigation revealed a tripped breaker and a blown fuse; the Duty Electrician resolved these problems and NuMI resumed taking beam. Unfortunately, the breaker began tripping again several minutes later. Tuesday Nov. 23 The midnight shift began with Operations monitoring store 8292 (128.49E30) and the stash (155E10), with Pbar stacking (14E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, at 7:19 AM, store 8292 quenched at D4. Operators investigated and found a failed Heater Firing Unit (HFU). The recovery of the cryo system lasted through to the end of the shift. Operators reported several more NuMI exhaust fan trips; the Duty Electrician adjusted the breaker trip point to get the system to stay up and running. Operations reported that a MI RF station (MIRF1) tripped off multiple times on gap envelope faults; Operations lowered the NuMI beam intensity, which reduced the frequency of these trips. Operations established beam to T-992 at 4 AM. During the day shift, Operations reported that the TeV was still in quench recovery mode. EE Support personnel replaced the offending HFU at 8:23 AM. Operations prepared CDF and D0 for collision hall accesses. The Cryo System Group conducted wet engine maintenance around the TeV ring. The NuMI exhaust fan tripped off again; FESS personnel discovered the trip was due to insulation wearing off one of the power cables; they replaced the cable at 10:35 AM. Operators replaced a ramp controller card for a quadrupole magnet (E:QF112). RF technicians replaced the modulator for MIRF1. Operators put the TeV through a dry squeeze and a wet squeeze. Operations put the TeV into shot setup at 1:30 PM, but put the shot on hold due to network communications issues. Operations put the TeV back into shot setup at 2:41 PM. Operations reported problems with the TeV SBD during shot setup. Operations reported that MTA was given permission for a special run, which was underway by the end of the shift. During the evening shift, Operations established store 8295 at 4:06 PM with an AIL of 301E30. Operations disabled the T-992 beam at 10 PM; this marked the end the experiment’s MTest run. Operators conditioned the MI gap cleaning kickers. Operators reported having communications issues with a Pbar router (PLC03); these issues resolved themselves as Pbar personnel investigated. Wednesday Nov. 24 The midnight shift began with Operations monitoring store 8295 (107.67E30) and the stash (201.2E10), with Pbar stacking (29E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 8295 at 7:15 PM and put the TeV into shot setup. The Plans for Wednesday and the long Weekend The plans for today and through the Weekend are to stack, stash, and store. Have a Happy Thanksgiving! |
|
More Information Comments and Suggestions |
| last modified 11/24/2010 email Fermilab |
| Security, Privacy, Legal | |